OpenStudy (sup_open~study):

Which measures would be the side lengths of a right triangle? A. 2 cm, 5 cm, 7 cm B. 9 cm, 14 cm, 17 cm C. 7 cm, 17 cm, 24 cm D. 5 cm, 12 cm, 13 cm

Parth (parthkohli):

Check it with the pythagorean theorem: \(\Large \color{MidnightBlue}{\Rightarrow a^2 + b^2 = c^2 }\) Note that the largest side would be c(hypotenuse as it is called)

OpenStudy (sup_open~study):

ok so i work all of them out how do i find out with on is the answer

Parth (parthkohli):

Try to check all if they satisfy the equation. I'll do one for you. \(\Large \color{MidnightBlue}{\Rightarrow 2^2 + 5^2 = 7^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow 4 + 25 \ne 49 }\) First is not the answer.

OpenStudy (sup_open~study):

what do you mean by satisfy the equation ???

Parth (parthkohli):

@ipm1988 Don't give the answer please, let the asker work on it.

Parth (parthkohli):

If the left hand side is equal to right hand side, then they may be the lengths of a right angled triangle.

OpenStudy (anonymous):

if it satisfy pythagorean theorem therefore it is a right triangle where c is the longest and a and b is the legs c^2=a^2 +b^2

OpenStudy (amorfide):

basically if A²+B² is equal to C² you have a right angled triangle

OpenStudy (sup_open~study):

o ok thanks guys

OpenStudy (anonymous):

@Sup_open~study D is the answer \[5^{2}+12^{2}\] = 169 sqrt(169)= 13

OpenStudy (sup_open~study):

yay i got it right thanks :)